# Learning to Compress Contexts for Efficient Knowledge-based Visual Question Answering

### v1

@@ Methods
1. RACC can utilize different types of knowledge sources to aid its efficient retrieval-augmented generation process. We evaluate RACC with two different knowledge sources, i.e. WIT and GS, which represent multimodal documents and textual documents. These two types of knowledge sources are quite common and important in practical applications, making them of significant research importance.
2. RACC can leverage any off-the-shelf multimodal retriever for retrieval, and our proposed RGCA strategy enables the retrieval results from high-performance retrievers to effectively guide the aggregation of the compressed prompts of retrieved documents. Therefore, RACC can benefit from advancements in multimodal retrieval technology.
3. RACC can be applied to any off-the-shelf MLLMs. In the setup of RACC-homo, the hyperMLLM, and the baseMLLM are identical, which means that the MLLM learns to compress contexts for itself. For RACC-hetero, the hyperMLLM and the baseMLLM differ in either structure or weights. We conduct experiments under this setup and present the results in Table 6. RACC-hetero also performs well across different baseMLLMs. The setup of RACC-hetero is also of practical significance: When it is not feasible to directly fine-tune the baseMLLM due to resource constraints, our framework can still work by adopting a much smaller hyperMLLM to help produce P-Tuningv2 modulations to adapt the frozen baseMLLM.
@@
BLIP2-Vicuna7B 60.67
InstructBLIP-Vicuna7B 62.91
@@
Conclusion
In this paper, we propose a KB-VQA framework named RACC, i.e. Retrieval-Augmented MLLMs with Compressed Contexts, which achieves efficient inference by learning to compress, aggregate and leverage retrieved contexts. The contributions of this paper can be summarized as follows:
1. RACC achieves state-of-the-art (SOTA) performance at a very low cost on the challenging OK-VQA datasets.
@@ Conclusion
3. RACC demonstrates broad applicability, as experiments show that it is applicable to different MLLMs and various kinds of external knowledge sources.
@@
With the rapid development of RAG technology and MLLMs, RAG on MLLMs will certainly attract increasing research attention. In the RAG of MLLMs, multimodal documents are very common and easily accessible in practical applications, which is of significant research importance. More importantly, we believe that inference latency is a key concern in practical applications, which has often been overlooked in previous KB-VQA works. For the above reasons, we conducted the research described in this paper and hope to provide some inspiration for future work.

### v2

@@ 4.6 Broad Applicability of RACC
1. RACC can utilize different types of knowledge sources to aid its efficient RAG process. We evaluate RACC with two knowledge sources, i.e. WIT and GS, which represent multimodal documents and textual documents.
2. RACC can leverage any off-the-shelf multimodal retriever for retrieval, and our proposed RGCA method enables RACC to benefit from advancements in multimodal retrieval technology.
3. RACC can be applied to any off-the-shelf MLLMs. We further conduct experiments under the setup of RACC-hetero and present the results in Table 6. Experiments show that RACC-hetero performs well across different baseMLLMs. The setup of RACC-hetero is also of practical significance: When it is not feasible to directly fine-tune the baseMLLM due to resource constraints, our framework can still work by adopting a much smaller hyperMLLM to adapt the larger frozen baseMLLM.
@@
BLIP2-Vicuna7B 61.65
InstructBLIP-Vicuna7B 63.92
@@
5 Conclusion
In this paper, we propose Retrieval-Augmented MLLMs with Compressed Contexts (RACC). RACC has achieved the following accomplishments in the area of KB-VQA:
1. RACC achieves competitive performance at a very low cost on challenging KB-VQA datasets.
@@ 5 Conclusion
3. RACC is applicable to different MLLMs and various kinds of external knowledge sources.
@@
With the rapid development of RAG technology and MLLMs, we believe that inference latency is a key concern in practical applications, which has often been overlooked in previous KB-VQA works. We hope our research will provide some inspiration for future work in this field.
Appendix A RAVQA-v2 reproduction
Since RAVQA-v2 does not investigate the effects of multimodal documents in its retrieval-agumented generation process, we replicate RAVQA-v2 using our crafted WIT knowledge source based on two types of MLLMs.
The experimental results are shown in Table 7. Specifically, we follow all the experiment setup mentioned by Lin et al. (2024a), such as the prompt template ”Question: {} Knowledge: {} Answer:”, where the Knowledge section includes both text tokens and image tokens transformed by Qformer. It is evident that the WIT knowledge source can, to some extent, provide the necessary knowledge for the OK-VQA dataset.
Additionally, we can observe that when using WIT as the knowledge source, the replication results of RAVQA-v2 across different MLLMs indicate that omitting the image information leads to better fine-tuning performance.
We believe this may be due to the following two reasons: (1) The inclusion of images directly increases the number of tokens in the knowledge section, making it more challenging for the MLLM to identify key information from the provided knowledge. (2) The use of images likely introduces redundant information that is unrelated to the document content, as images themselves contain a wealth of information. Furthermore, incorporating images also increases the time required for both training and inference based on RAVQA-v2.
Overall, although RAVQA-v2 theoretically supports multimodal document-based knowledge sources, experiments show that its performance with multimodal documents is not satisfactory.
MLLM Knowledge Source VQA Accuracy (%)
BLIP2 No Knowledge 54.10
WIT 56.26
WIT (Text Only) 56.44
InstructBLIP No Knowledge 57.32
WIT 58.77
WIT (w/o Image) 59.12
Table 7: The performance of two types of MLLMs after finetuned on OK-VQA with the WIT knowledge source, based on the RAVQA-v2 framework.
Appendix B Complete Ablation Studies
We identify four key issues in training RACC and further propose four methods to solve the issues. Extensive ablation studies are conducted to verify the effectiveness the proposed four methods.
We conduct experiments on two types of MLLMs, i.e. BLIP2-FlanT5XL and InstructBLIP-FlanT5XL. Due to page constraints of the paper content, we only present the ablation results for InstructBLIP-FlanT5XL in Table 8. We adopt RACC-homo in the experiments of ablation studies, where the hyperMLLM and baseMLLM are both initialized from the models listed in the ”MLLM” column of Table 8.
For RACC-homo based on BLIP2-FlanT5XL, comparing lines 2 and 3, as well as lines 7 and 8, we can observe that the PIPE method brings improvements of 0.28% and 0.44% under different settings. The PIPE method is also effective for the InstructBLIP-based RACC-homo, leading to performance gains of 0.31% and 0.42% as shown by lines 10 and 11, as well as lines 15 and 16.
From the differences between lines 3 and 4, as well as lines 11 and 12 in Table 5, we can observe the effectiveness of the DCSE method, bringing improvements of 0.55% and 0.77% respectively. The RGCA method, on the other hand, results in improvements of 0.28% and 0.37%, as shown in lines 3 and 5, as well as lines 11 and 13 in Table 5.
Last but not least, the performance differences between lines 6 and 8 and lines 14 and 16 in Table 5 show that the PRDB method leads to performance gains of 0.39% and 0.54% based on the two types of MLLMs.
We also plot the VQA accuracy on the validation set during the training process of the framework using different methods, as shown in Figure 2.
Figure 2: The VQA accuracy on the validation set during the training process of the framework using different methods.
MLLM No. PIPE DCSE RGCA PRDB VQA Accuracy (%)
BLIP2-FlanT5XL 1 53.77
2 54.20(+0.43)
3 54.48 (+0.71)
4 55.03 (+1.26)
5 54.76 (+0.99)
6 54.81 (+1.04)
7 54.76 (+0.99)
8 55.20 (+1.43)
InstructBLIP-FlanT5XL 9 57.60
10 58.18 (+0.58)
11 58.49 (+0.89)
12 59.26 (+1.66)
13 58.86 (+1.26)
14 58.95 (+1.35)
15 59.07 (+1.47)
16 59.49 (+1.89)
Table 8: The results of ablation studies on the design of our proposed framework. The GS knowledge source is adopted here. The ablation experiments are conducted based on RACC-homo with the above two types of MLLMs.
Appendix C Details of Different Knowledge Sources
In this section, we further introduce more details of the two knowledge sources, i.e. GS and WIT.
First of all, the word count in knowledge source documents directly affects the efficiency during training and inference in RAG applications. Therefore, we first present some statistics of the document word counts for these two knowledge sources in Table 9.
We employ FLMR and PREFLMR to retrieve documents from these knowledge sources. FLMR has been trained on OK-VQA and GS, thus it has a strong capability to retrieve relevant documents from the GS knowledge source. PREFLMR is an upgraded version of FLMR, equipped with powerful capabilities for retrieving multimodal documents that FLMR lacks. The detailed retrieval results for OK-VQA with different knowledge sources are shown in Table 10.
Firstly, from Table 10, we can tell that there is obviously a number of irrelevant documents among the retrieved documents in different knowledge sources.
Second, Table 11 shows that the quality of the GS knowledge base is superior to that of the WIT knowledge base. The PPRecall@1 metric for the GS knowledge base reaches 63.59%, indicating that the document with the highest confidence retrieved has a high probability of containing the answer to the question. The PRRecall@5 metric even reaches 88.32%. These data highlight the advantages of GS as a knowledge source specifically designed for OK-VQA.
Knowledge Source Metric Statistic
GS Mean 59.76
Std 46.05
Min 10
25% 29.0
50% 45.0
75% 75.0
Max 521
WIT Mean 155.71
Std 78.86
Min 24
25% 89
50% 149
75% 214
Max 1122
Table 9: The statistical data on document word counts of different knowledge sources. The ”25%” column indicates that 25% of the documents have a word count below this value, and similarly for other percentiles.
Knowledge Source Retriever K PRRecall@K (%)
Train Val
GS FLMR 1 71.56 63.59
2 83.77 77.07
3 88.63 82.85
4 91.03 86.41
5 92.82 88.32
10 95.99 93.63
WIT PREFLER 1 37.09 35.85
2 48.26 46.76
3 54.89 53.27
4 59.37 57.67
5 62.89 60.93
10 72.16 70.41
Table 10: The detailed retrieval results of different knowledge sources. PRRecall@K measures whether the retrieved $K$ documents contain at least one pseudo-relevant document..
Appendix D Investigation of hyperparameter $K$
The hyperparameter $K$ represents the number of retrieved documents used for each image-question pair. The value of $K$ is typically set to 5 in this paper, consistent with RAVQA-v2. We also investigated the impact of different values of $K$ across various knowledge sources and the results are shown in Table 11.
On the GS knowledge base, when $K=1$, RACC-homo achieved a performance of 59.65%, surpassing the 59.49% performance when $K=5$. This may be due to the high quality of the GS knowledge base, where the first retrieved document is often likely to contain the information needed to answer the question. In such cases, setting $K$ to a larger value might directly introduce documents with weaker relevance, thereby affecting the performance of the whole framework.
On the WIT knowledge source, we conduct experiments with and without images. In both settings, a larger value of $K$ results in better performance. This may be because there are fewer documents related to the question in the WIT knowledge source, so a larger $K$ allows RACC to extract useful information from more documents. Furthermore, for the same value of $K$, using images results in better performance in both settings, indicating that RACC can obtain valuable information from images. Compared to the results in Table 7, we can conclude that RACC has an advantage over RAVQA2 when utilizing multimodal document knowledge sources.
Finally, we can conclude from Table 11 that the RACC framework achieves excellent performance across various knowledge sources and different settings of $K$.
knowledge source K PRRecall@K (%) VQA Accuracy (%)
WIT 1 35.85 59.11
3 53.27 59.16
5 60.93 59.17
WIT (Text Only) 1 35.85 58.97
3 53.27 59.01
5 60.93 59.04
GS 1 63.59 59.65
3 82.85 59.35
5 88.32 59.49
Table 11: RACC-homo’s experimental results with varying hyperparameter $K$ across different knowledge sources. The MLLM used here is InstructBLIP-FlanT5XL.
Appendix E Pseudo Code of RGCA method
Our proposed RGCA method is applied during the forward computation process of the cross-attention mechanism, aiming to ensure that compressed prompts corresponding to documents with higher retrieval scores receive more attention in the retrieval-guided cross-attention blocks. The pseudo code for the forward function of the retrieval-guided cross-attention block is as follows:
Listing 1 Pseudo code of RGCA
1def forward(self, x, context, r_scores):
2 # projecting inputs into q, k, v
3 # transform q, k, v into multi-head forms
4 # expanding the retrieval scores to a certain shape
5 r_scores = repeat(r_scores, ’b x y -> (b h) (x m) (y n)’, h=self.heads, m=x.shape[1], n=context.shape[1])
6
7 # calculating the similarities between queries and keys
8 sim = einsum(’b i d, b j d -> b i j’, q, k) * self.scale
9 sim = sim * r_scores
10 attn = sim.softmax(dim=-1)
11
12 # calculating the outputs